# The Orientalists’ Stance Towards Arabic Sciences (Especially Arabic Astronomy)

### v1

@@ Abstract
In this paper, we highlight the influence of Arab/Islamic civilization in the field of the history of astronomy on European historians. We also aim to elucidate the stance of Orientalists toward the study of Arab sciences and to clarify their orientations, with a particular focus on astronomy, while revealing the significant role played by Arab scholars in this domain and the impact of their contributions—especially astronomical tables (zīj)—on Western astronomers. Furthermore, we have clarified the mechanisms of transmission of Arab sciences, particularly astronomy, from Arab scholars to Western scholars, and the role of Arab astronomers in Western civilization. In addition, we address the contributions of Arab scholars to the development of astronomy and the perspective of Orientalists, particularly David King, regarding this matter. We also underscore the importance of Orientalists’ works in analyzing Arab/Islamic scholarly output, identifying its influence on the West in the field of astronomy, and demonstrating how Western scholars benefited from translations of Arabic books in this discipline. In this paper, we adopt the historical retrieval methodology, by referencing previously documented astronomical information and contributions, with an emphasis on the processes of transmission of these sciences from the Arabs to the West.
@@ 1 Introduction
The influence of Arab/Islamic civilization on the Western world was profoundly significant, as numerous astronomical works—particularly Arabic astronomical tables (zījes)—were transmitted to them. The most compelling evidence lies in the acknowledgment by many Orientalists of the monumental role played by Arab/Islamic culture in enriching European thought. Astronomy owes a substantial portion of its advancement to the Arabs, whose contributions played a pivotal role in its development, preservation, and transmission to the West through translation, refinement, correction via precise observation and verification, and subsequent conveyance.

### v2

@@ Abstract
In this paper, we highlight the influence of Arab/Islamic civilization in the field of the history of astronomy on European historians. We also aim to elucidate the stance of Orientalists toward the study of Arab sciences and to clarify their orientations, with a particular focus on astronomy, while revealing the significant role played by Arab scholars in this domain and the impact of their contributions—especially astronomical tables (zij)—on Western astronomers. Furthermore, we have clarified the mechanisms of transmission of Arab sciences, particularly astronomy, from Arab scholars to Western scholars, and the role of Arab astronomers in Western civilization. In addition, we address the contributions of Arab scholars to the development of astronomy and the perspective of Orientalists, particularly David King, regarding this matter. We also underscore the importance of Orientalists’ works in analyzing Arab/Islamic scholarly output, identifying its influence on the West in the field of astronomy, and demonstrating how Western scholars benefited from translations of Arabic books in this discipline. In this paper, we adopt the historical retrieval methodology, by referencing previously documented astronomical information and contributions, with an emphasis on the processes of transmission of these sciences from the Arabs to the West.
@@ 1 Introduction
The influence of Arab/Islamic civilization on the Western world was profoundly significant, as numerous astronomical works—particularly Arabic astronomical tables (zijes)—were transmitted to them. The most compelling evidence lies in the acknowledgment by many Orientalists of the monumental role played by Arab/Islamic culture in enriching European thought. Astronomy owes a substantial portion of its advancement to the Arabs, whose contributions played a pivotal role in its development, preservation, and transmission to the West through translation, refinement, correction via precise observation and verification, and subsequent conveyance.